OpenStudy (lifeisadangerousgame):

Radicals ._. @ganeshie8 Am I doing this right?

OpenStudy (lifeisadangerousgame):

First I had the problem then I put \[\sqrt{8 \times 6 \times y \times y}\] Is that right so far?

ganeshie8 (ganeshie8):

nope, not ok cuz you're doing it reverse

ganeshie8 (ganeshie8):

\(\large \sqrt{48y^2}\)

ganeshie8 (ganeshie8):

convert the radical to exponent

ganeshie8 (ganeshie8):

\(\large \sqrt{48y^2}\) \(\large (48y^2)^{\frac{1}{2}}\)

ganeshie8 (ganeshie8):

\(\large \sqrt{48y^2}\) \(\large (48y^2)^{\frac{1}{2}}\) \(\large (48)^{\frac{1}{2}} (y^2)^{\frac{1}{2}}\)

ganeshie8 (ganeshie8):

\(\large \sqrt{48y^2}\) \(\large (48y^2)^{\frac{1}{2}}\) \(\large (48)^{\frac{1}{2}} (y^2)^{\frac{1}{2}}\) \(\large (48)^{\frac{1}{2}}y\)

ganeshie8 (ganeshie8):

fine so far ?

OpenStudy (lifeisadangerousgame):

Fine so far

ganeshie8 (ganeshie8):

next, use this :- 48 = 16x3

ganeshie8 (ganeshie8):

\(\large \sqrt{48y^2}\) \(\large (48y^2)^{\frac{1}{2}}\) \(\large (48)^{\frac{1}{2}} (y^2)^{\frac{1}{2}}\) \(\large (48)^{\frac{1}{2}}y\) \(\large (16 \times 3)^{\frac{1}{2}}y\) \(\large (16)^{\frac{1}{2}} (3)^{\frac{1}{2}} y\)

ganeshie8 (ganeshie8):

\(\large \sqrt{48y^2}\) \(\large (48y^2)^{\frac{1}{2}}\) \(\large (48)^{\frac{1}{2}} (y^2)^{\frac{1}{2}}\) \(\large (48)^{\frac{1}{2}}y\) \(\large (16 \times 3)^{\frac{1}{2}}y\) \(\large (16)^{\frac{1}{2}} (3)^{\frac{1}{2}} y\) \(\large (4^2)^{\frac{1}{2}} (3)^{\frac{1}{2}} y\)

ganeshie8 (ganeshie8):

\(\large \sqrt{48y^2}\) \(\large (48y^2)^{\frac{1}{2}}\) \(\large (48)^{\frac{1}{2}} (y^2)^{\frac{1}{2}}\) \(\large (48)^{\frac{1}{2}}y\) \(\large (16 \times 3)^{\frac{1}{2}}y\) \(\large (16)^{\frac{1}{2}} (3)^{\frac{1}{2}} y\) \(\large (4^2)^{\frac{1}{2}} (3)^{\frac{1}{2}} y\) \(\large 4 (3)^{\frac{1}{2}} y\)

ganeshie8 (ganeshie8):

\(\large \sqrt{48y^2}\) \(\large (48y^2)^{\frac{1}{2}}\) \(\large (48)^{\frac{1}{2}} (y^2)^{\frac{1}{2}}\) \(\large (48)^{\frac{1}{2}}y\) \(\large (16 \times 3)^{\frac{1}{2}}y\) \(\large (16)^{\frac{1}{2}} (3)^{\frac{1}{2}} y\) \(\large (4^2)^{\frac{1}{2}} (3)^{\frac{1}{2}} y\) \(\large 4 \sqrt{3} y\) \(\large 4y \sqrt{3}\)

ganeshie8 (ganeshie8):

Done. see if ti makes soem sense :)

OpenStudy (lifeisadangerousgame):

like it makes sense how you got the answer, I can see how you got it and the steps, but I still don't understand how I would do it if I had to do this on my own

ganeshie8 (ganeshie8):

ikr.... practice few problems, until u feel confident. there is no other way !

OpenStudy (lifeisadangerousgame):

Definitely, thanks Ganeshie!

ganeshie8 (ganeshie8):

np :) remember this :- when u pull 4^2 out from a radical, it becomes 4

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):